# MSDNet: Multi-Scale Decoder for Few-Shot Semantic Segmentation via Transformer-Guided Prototyping

### v1

@@ IV-E Cross-dataset task
Our experimental results, as detailed in Table III, demonstrate the superior performance of our proposed method compared to existing SOTA approaches under both 1-shot and 5-shot evaluation scenarios. This underscores the robustness and effectiveness of our few-shot segmentation framework in handling cross-dataset challenges and domain shifts.
IV-F Ablation Study
For the ablation study, we conduct experiments on the $COCO-20^{i}$ dataset using the ResNet50 backbone in 1-shot scenario. Our first aim is to investigate the individual impact of various components on the segmentation performance. Table IV show the impact of each component of the proposed method.
The first row of Table IV represents the performance of the baseline model, consisting solely of the backbone architecture and support prototype mechanism. Subsequent rows introduce additional components incrementally, including the CMGM, STD, and multi-scale decoder.
TABLE IV: The Impact of Each Component on Segmentation Performance in the $COCO-20^{i}$ Dataset
@@ IV-F Ablation Study
$\checkmark$ 30.1 34.2 33.4 33.8 32.9 59.7
$\checkmark$ $\checkmark$ 31.5 35.9 34.8 34.2 34.1 60.8
$\checkmark$ $\checkmark$ $\checkmark$ 43.0 45.2 43.1 41.4 43.2 67.6
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 43.7 49.1 46.9 46.2 46.5 70.4
As shown in Table IV, each component contributes to an improvement in segmentation performance, with the Multi Scale Decoder showcasing the most substantial impact. The progressive integration of these components results in a notable enhancement in mIoU scores across all folds, underscoring their significance in refining segmentation masks and capturing contextual information effectively.
Furthermore, in Figure 5, we present the qualitative results obtained by incorporating each component into the baseline model on the $COCO-20^{i}$ dataset. As illustrated in Figure 5, the addition of each component leads to noticeable improvements in the segmentation results. Particularly, the integration of the multi-scale decoder component demonstrates significant enhancement in segmentation accuracy.
To further explore the influence of the architecture within the multi-scale decoder, we conducted an ablation study varying the number of residual blocks in each stage. Figure 6 provides an overview of the Multi-Scale Decoder with different numbers of residual blocks in each stage. The experiment involved evaluating the segmentation performance on the $COCO-20^{i}$ dataset using the ResNet50 backbone in a 1-shot scenario. As depicted in Table V, we examined configurations ranging from one to four residual blocks per stage. Interestingly, the results revealed that the optimal segmentation performance was achieved with three residual blocks in each stage. This finding suggests that an appropriate balance in the depth of the decoder architecture plays a crucial role in enhancing segmentation accuracy. Too few blocks may limit the model’s capacity to capture intricate features, while an excessive number of blocks could lead to overfitting or computational inefficiency. Therefore, our results underscore the importance of carefully tuning the architecture parameters to achieve optimal performance in few-shot segmentation tasks.
TABLE V: The Impact of number of residual blocks in each stage of Multi Scale Decoder on Segmentation Performance in the $COCO-20^{i}$ Dataset
@@
V Conclusion
In conclusion, our proposed few-shot segmentation framework, leveraging a combination of components including a shared pretrained backbone, support prototype mechanism, CMGM, STD, and multi-scale decoder, has demonstrated remarkable efficacy in achieving SOTA performance on both $PASCAL-5^{i}$ and $COCO-20^{i}$ datasets. Through extensive experimentation and ablation studies, we have highlighted the critical contributions of each component, particularly emphasizing the significant impact of the multi-scale decoder in enhancing segmentation accuracy while maintaining computational efficiency. Looking ahead, further investigation into the dynamic adaptation of prototype representations and the exploration of additional attention mechanisms could offer avenues for improving the adaptability and robustness of our method across diverse datasets and scenarios. Additionally, exploring semi-supervised learning paradigms could enhance the generalization capability of our framework, enabling effective segmentation in scenarios with limited labeled data. These avenues for future work hold promise for advancing the effectiveness and applicability of few-shot segmentation methods in real-world scenarios.

### v3

@@ 4.5 Cross-dataset task
Our experimental results, as detailed in Table 3, demonstrate the superior performance of our proposed method compared to existing SOTA approaches under both 1-shot and 5-shot evaluation scenarios. This underscores the robustness and effectiveness of our few-shot segmentation framework in handling cross-dataset challenges and domain shifts.
4.6 Ablation Study
(a) Qualitative comparison of component effects on $COCO\text{-}20^{i}$ dataset in 1-shot scenario
(b) Qualitative comparison of component effects on $Pascal\text{-}5^{i}$ dataset in 1-shot scenario
Figure 4: Qualitative comparison of component effects in 1-shot scenario for (a) $COCO\text{-}20^{i}$ and (b) $Pascal\text{-}5^{i}$ datasets.
To evaluate the contribution of each proposed component, we perform an ablation study on the $COCO\text{-}20^{i}$ dataset using the ResNet50 backbone under the 1-shot setting. The results are summarized in Table 4.
The first row of Table 4 shows the baseline performance, which includes only the backbone and the support prototype mechanism. In the following rows, we incrementally introduce each component—namely, CMGM, STD, and the Multi-Scale Decoder—to analyze their individual and combined effects on segmentation performance.
Table 4: The Impact of Each Component on Segmentation Performance in the $COCO-20^{i}$ Dataset
@@ 4.6 Ablation Study
[…]
As shown in Table 4, each component contributes to an improvement in segmentation performance, with the Multi Scale Decoder showcasing the most substantial impact. The progressive integration of these components results in a notable enhancement in mIoU scores across all folds, underscoring their significance in refining segmentation masks and capturing contextual information effectively.
Table 5: The Impact of number of residual blocks in each stage of Multi Scale Decoder on Segmentation Performance in the $COCO-20^{i}$ Dataset
@@
Furthermore, in Figure 4, we present a qualitative comparison illustrating the effect of progressively adding each proposed component to the baseline model on two benchmark datasets: $COCO\text{-}20^{i}$ and $Pascal\text{-}5^{i}$, both under the 1-shot setting. Specifically, Figure 4-4(a) shows the results on the $COCO\text{-}20^{i}$ dataset, while Figure 4-4(b) displays the corresponding outcomes on the $Pascal\text{-}5^{i}$ dataset. As observed in both subfigures, the incorporation of each component consistently leads to noticeable improvements in segmentation quality. In particular, the introduction of the multi-scale decoder contributes significantly to capturing fine-grained details and enhancing object boundaries, thereby demonstrating its effectiveness in improving the overall segmentation performance across diverse image domains.
To further explore the influence of the architecture within the multi-scale decoder, we conducted an ablation study varying the number of residual blocks in each stage. Figure 5 provides an overview of the Multi-Scale Decoder with different numbers of residual blocks in each stage. The experiment involved evaluating the segmentation performance on the $COCO-20^{i}$ dataset using the ResNet50 backbone in a 1-shot scenario. As depicted in Table 5, we examined configurations ranging from one to four residual blocks per stage. Interestingly, the results revealed that the optimal segmentation performance was achieved with three residual blocks in each stage. This finding suggests that an appropriate balance in the depth of the decoder architecture plays a crucial role in enhancing segmentation accuracy. Too few blocks may limit the model’s capacity to capture intricate features, while an excessive number of blocks could lead to overfitting or computational inefficiency. This experiment also reflects our effort to maintain a lightweight design without compromising performance. The resulting model achieves a low parameter count (1.5M) not through arbitrary reduction, but through deliberate architectural tuning, ensuring both effectiveness and efficiency. Therefore, our results underscore the importance of carefully tuning the architecture parameters to achieve optimal performance in few-shot segmentation tasks.
5 Conclusion
In conclusion, our proposed few-shot segmentation framework, leveraging a combination of components including a shared pretrained backbone, support prototype mechanism, CMGM, STD, and multi-scale decoder, has demonstrated remarkable efficacy in achieving SOTA performance on both $PASCAL-5^{i}$ and $COCO-20^{i}$ datasets. Through extensive experimentation and ablation studies, we have highlighted the critical contributions of each component, particularly emphasizing the significant impact of the multi-scale decoder in enhancing segmentation accuracy while maintaining computational efficiency. While our method shows strong performance, it is not without limitations. First, the use of a fixed support prototype may oversimplify the representation of intra-class variance in some complex categories. This can lead to reduced accuracy when the support and query images differ significantly in appearance. Second, although our model is lightweight, the presence of dual decoder modules (STD and multi-scale decoder) introduces additional inference time compared to simpler architectures. Lastly, the current architecture is tailored for single-class segmentation per episode; extending it to multi-class few-shot scenarios would require further adaptation and optimization. Looking ahead, further investigation into the dynamic adaptation of prototype representations and the exploration of additional attention mechanisms could offer avenues for improving the adaptability and robustness of our method across diverse datasets and scenarios. Additionally, exploring semi-supervised learning paradigms could enhance the generalization capability of our framework, enabling effective segmentation in scenarios with limited labeled data. These avenues for future work hold promise for advancing the effectiveness and applicability of few-shot segmentation methods in real-world scenarios.